OpenStudy (aonz):

Halp please. Medal will be awarded Express in terms of log 2 (3) and log 2 (5) log 2 (5/8)

OpenStudy (aonz):

@ash2326 @jim_thompson5910 Help please @whpalmer4

OpenStudy (aonz):

@hartnn halp please :)

OpenStudy (aonz):

they are all base 2 btw

hartnn (hartnn):

not sure on what exactly the question is asking, but you can use : \(\large \log _a(b/c)= \log_a b -\log_ac\) so, what you'll get for \(\log_2 (5/8)=...?\)

OpenStudy (aonz):

log2(5) - log2(8)

hartnn (hartnn):

yes, how can you write 8 as the power of 2 ?

OpenStudy (aonz):

2^3 ahhh :) and 2 and 2 cuts off

OpenStudy (aonz):

so log2(5) -3 is final answer?

hartnn (hartnn):

yes, because \(\large \log_22=1\)